OpenStudy (anonymous):

Hi! I'm in Pre-Calculus and we are working with functions (domain and range etc.) I'm stuck on graphing f(x)=int(2x). Could you help? :)

OpenStudy (loser66):

Precal and int? no way , right? or f( x) = sin (2x)?

OpenStudy (anonymous):

Well it is chapter 2 and we are doing int....i guess. But I don't know what that means. Are we like really behind? :(

OpenStudy (loser66):

to me, int is integral which I was taught near end of calculus1 course. At precal, I just studied sin , cos, .... (trigs) only

OpenStudy (anonymous):

oh :) my teacher said something about integrals and he tried to explain it but it made no sense because i've never seen it before. I took an Algebra II/Trig class last year and we covered a lot of pre-calc in it, but never integrals.

OpenStudy (loser66):

you didn't study int yet, how can you sketch that graph? Although it 's quite simple, but if you don't know how to take int (2x) dx , I don't know how to say.....

OpenStudy (loser66):

ok, pretend as if you know it. hehehe. Pretend as if what I say make a perfect sense to you ( I am a liar!!) \(\int 2xdx = x^2 + C\) so f(x)= \(x^2 +C\) are the family graph of parabola

OpenStudy (anonymous):

ok......:) I'm following you (very slowly)

OpenStudy (loser66):

whose vertex is on the y axis. if